A black mamba travels 20 miles per hour for 1/4 mile. How many seconds does it take to travel the 1/4 mile?

Answer by Alan3354(69443)

A black mamba travels 20 miles per hour for 1/4 mile. How many seconds does it take to travel the 1/4 mile?
------------
d = r*t
t = d/r = 0.25/20
t = 0.0125 hours
---------
0.0125 hr * 60 min/hr * 60 sec/min = 45 seconds
=================
Notice how including the units clarifies it.